OpenStudy (rollfaceblah):

Help please?? The leg of a right triangle is 3 units and the hypotenuse is 7 units. What is the length, in units, of the other leg of the triangle?

OpenStudy (anonymous):

do you have an idea of how to do this?

OpenStudy (rollfaceblah):

I really don't its a homework question

OpenStudy (anonymous):

\[a ^{2}+b ^{2}=c ^{2}\]

OpenStudy (anonymous):

leg1=A leg2=b hypot=c

OpenStudy (rollfaceblah):

so 3^2 + 7^2= c^2? c^2 = 58 c= 7.615??